OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)?

OpenStudy (anonymous):

to find the slope you need to use the slope formula m = (y1-y2)/(x1-x2) using the points given (3, -1) as (x1, y1) and (-2, -5) as (x2, y2) you will get m = (-1 - (-5)) / (3 - (-2)) = (-1 + 5)/ (3 + 2) = (4/5) m = 4/5

OpenStudy (anonymous):

thnx